# A Knowledge Graph-based Retrieval-Augmented Generation Framework for Algorithm Selection in the Facility Layout Problem

### v1

@@ 3.1 Evaluation methodology
      Reasoning quality: A qualitative score from 1 (Poor) to 5 (Excellent) assessing the extent to which the LLM’s explanation was detailed, rich, and verifiably based on the evidence provided. The reasoning quality for both methods was evaluated by a separate instance of the Google Gemini 2.5 Flash LLM, which was prompted to act as an impartial judge and score the outputs based on the following criteria.
      1. (a)
        1 (Poor): Irrelevant or ambiguous explanation.
      2. (b)
        2 (Weak): Plausible but not specific.
      3. (c)
        3 (Acceptable): Reasonable but unsupported by cited evidence.
      4. (d)
        4 (Good): Applies to some numbers, but the analysis is superficial.
      5. (e)
        5 (Excellent): Combines several correct, cited evidence points to construct a firm, evidence-based conclusion.
3.2 Results and discussion
Table 4 compares outcomes of the proposed and baseline model with the ground truth and the data-driven ranking based on their explainability.
@@ 3.2 Results and discussion
3.2.1 Analysis of baseline method performance
While capable of reasoning, the baseline LLM’s recommendations achieve an overall accuracy of only 20% (1 out of 5). Its failures highlight the limitations of reasoning over structured tabular data. In Test cases 1 & 4, the baseline failed to recommend the ground truth algorithm because its reasoning was limited to a superficial raw text analysis. It could not perform the high-level frequency analysis (like the KG-RAG’s cluster analysis) needed to identify the most applied and successful methods. Moreover, it fails to interpolate the number of facilities when the exact number of facilities is not present in the dataset, as in test case 4. In Test case 2, the baseline’s recommendation of ACO-FBS for a multi-objective problem was factually incorrect. A query against the provided dataset would have revealed that ACO-FBS is only associated with Single-Objective problems. The baseline, lacking this explicit relationship, made a plausible but technically flawed suggestion, resulting in a reasoning score of 1.
Figure 5: Model recommendation rating analysis
3.2.2 Analysis of KG-RAG method performance
The KG-RAG method performed exceptionally well, achieving 100% recommendation accuracy and a perfect 5/5 reasoning score on all test cases. Its strength lies in its ability to synthesize multiple layers of evidence from the structured graph. In Test case 1, the KG-RAG model’s reasoning addressed specific problem instances and higher-level trends from the objective and scale clusters. This allowed it to correctly identify CRO-SL as the top contender, a conclusion the baseline missed. In Test case 2, the method’s ability to explicitly query for problems within the multi-objective cluster was critical. This ensured that the evidence provided to the LLM was highly relevant, leading to the correct recommendation of HSA. In Test case 5, even when the query did not contain an objective or constraints, the KG-RAG method provided a more nuanced and helpful recommendation. Analysing the broader context of 30-facility problems suggested both a simple heuristic (Construction Heuristic) and a more powerful metaheuristic (BRKGA-LP), giving the user a better understanding of the available trade-offs. This shows that the evaluation confirms that while a baseline LLM can reason from raw data to a limited extent, the KG-RAG model’s ability to leverage a structured, connected knowledge base allows it to perform an accurate analysis, yielding superior recommendations and data-driven explanations.
@@ 4 Conclusion
The paper introduced a new recommendation method based on a KG-RAG framework for FLP. This method successfully integrates the formalized reasoning strength of a Neo4j graph database with the advanced compilation strength of an LLM to generate explainable, high-quality recommendations.
The primary contribution of this paper is a multi-stage pipeline that integrates graph-based, vector-based, and cluster-based search approaches to construct a rich, evidence-supported context for every user query. The evaluation strongly confirmed the benefit of this approach. The proposed KG-RAG method significantly outperformed a baseline LLM with direct access to the raw data, achieving 100% recommendation accuracy against the ground truth, compared to 20% for the baseline. Furthermore, the quality of the data-driven reasoning was consistently rated as an average score of 5/5, far exceeding the baseline’s average score (3.4/5). This performance stems from the KG-RAG method’s ability to synthesize high-level trends from clusters with granular evidence from individual problem nodes, resulting in accurate, transparent, and reliable justifications. The risk of hallucination is substantially reduced by grounding the LLM in a domain-specific knowledge graph.
Despite these promising results, the proposed method has certain limitations. The method’s effectiveness fundamentally depends on the quality and scope of its underlying KG. It is a domain-specific tool for the FLP and recommendations for other domain optimization problems without a new, relevant knowledge base. Additionally, while the method includes a fallback mechanism for out-of-scope queries, for example, the problems with a number of facilities exceeding the KG’s included maximum data, its recommendations in these scenarios are based on broader categorical data rather than direct, comparable evidence, making them inherently less precise. Addressing this requires continuously expanding the knowledge base over time through the intelligent feedback loop. Furthermore, this paper prescribes a tangible and scalable way of building continuously updated methods. The collective feedback loop, which continuously feeds and aggregates new solutions populated by the users, maintains the knowledge graph updated over time. This makes the method a dynamic knowledge base that updates itself based on the user’s input. Future work will attempt to extend the framework to other challenging combinatorial optimization problems.

### v2

@@ 2.5.1 Evaluation methodology
      Reasoning quality: The reasoning quality for both methods was evaluated by a separate instance of the Google Gemini 1.5 Flash LLM, which was prompted to act as an impartial judge. It scored the outputs on a scale where 1 (Poor) indicated an irrelevant or ambiguous explanation, and 2 (Weak) denoted a plausible but non-specific argument. A score of 3 (Acceptable) was given for reasonable but unsupported claims, while 4 (Good) was assigned to explanations that used some data but where the analysis remained superficial. The highest score, 5 (Excellent), was reserved for responses that combined multiple, correct, and cited evidence points to construct a firm, evidence-based conclusion.
3 Results and discussion
The experimental evaluation focuses on representative Facility Layout Problem (FLP) cases that span variations in facility scale, objective structure, and geometric or operational constraints. These cases capture typical FLP challenges such as minimizing material-handling cost, balancing conflicting objectives, and satisfying constraints like non-overlap, boundary limits, area requirements, and aspect-ratio conditions. Evaluating the proposed method on this diverse set of instances allows us to assess its reasoning capability across problem types commonly encountered in facility-planning contexts. This section interprets the experimental results, compares the performance of the proposed KG-RAG method with the baseline LLM, and discusses the practical implications, failure modes, and operational boundaries revealed through the test cases. Where relevant, the discussion connects these observations to the retrieval components and learning mechanisms described in Section 2 to clarify why specific outcomes were obtained and how they extend to related FLP scenarios.
The comparative evaluation shows a clear advantage for the KG-RAG method: across the six test cases, it achieved an overall average reasoning score of approximately 4.7/5, while the baseline LLM, when provided with the raw CSV as its knowledge source, achieved an average reasoning score of about 3.3/5, as shown in Table 4 and Figure 5. These results indicate that the structured relations encoded in the KG materially improve instance-level matching and evidence provenance, whereas the baseline relies primarily on textual co-occurrence and pattern matching. The baseline’s principal error pattern was inferring algorithm suitability from lexical proximity in the CSV rather than from formally encoded relationships. For example, in test case 2 (a multi-objective instance), the baseline recommended ACO-FBS despite the dataset linking ACO-FBS only to single-objective problems. The KG-RAG method avoided this error by restricting retrieval to the multi-objective cluster and by providing the LLM solely with multi-objective precedents. The qualitative reasoning scores further highlight the practical difference between methods. KG-RAG outputs consistently contained traceable citations to problem nodes, method nodes, and cluster trends, which the judging LLM rated as strong evidence-based reasoning. In contrast, baseline outputs often presented plausible but weakly substantiated arguments. Taken together, these findings show that the KG-RAG approach substantially improves explainability and verifiability, even when the evaluation includes boundary-scale cases at the upper limit of KG coverage.
Figure 5: Model recommendation rating analysis
@@ 3 Results and discussion
6 No of facilities = 125, Objective = None, Constraints = None BRKGA-LP SDP 3 SDP 3
Examining each test case in more depth clarifies how the retrieval channels contributed to the observed outcomes. In the sixth test case (125 facilities), the system did not retrieve the ground-truth BRKGA-LP because the knowledge graph contains very limited evidence at this upper scale, causing both KG-RAG and the baseline to fall back to cluster-level patterns and converge on the same SDP recommendation with a lower reasoning score. In test case 1 (10 facilities, minimize material handling cost), the graph search retrieved several closely related precedents that supported CRO-SL as a proven algorithm for this problem size, while the cluster trends confirmed that BRKGA variants are frequently applied to similar small- and medium-scale instances. The combination allowed the system to recommend both an instance-backed algorithm and a strong alternative. In test case 2 (multi-objective), the restriction of retrieval to the explicitly defined multi-objective cluster prevented a category mismatch and enabled the correct selection of HSA. The baseline failed here by recommending ACO-FBS, which appeared in the dataset under partial feature matches but is only valid for single-objective cases. In test case 4 (40 facilities), exact precedents were absent in the KG. The KG-RAG method addressed this gap by interpolating across a ±25% facility range and incorporating evidence from the large-scale cluster. This fallback ensured a useful recommendation supported by aggregate patterns, even when instance-level evidence was sparse. Beyond these designed test cases, additional deployment scenarios highlight the broader utility of the framework. In dynamic reconfiguration problems, where objectives change as layouts evolve, the pipeline can requery objective and scale clusters to surface algorithms historically applied under similar shifts, providing actionable guidance without requiring new data collection. For highly constrained layouts involving adjacency or aspect-ratio conditions, the graph’s explicit links between problems and constraint-handling strategies make it possible to identify hybrid algorithms, such as metaheuristics combined with LP repair operators, that have succeeded in comparable contexts. Finally, for novice users or under-specified queries, the combined use of cluster-level trends and vector similarity led the system to propose both a conservative starting point (a construction heuristic) and a higher-performance alternative (a BRKGA variant), demonstrating how the retrieval channels interact to generate useful recommendations even when minimal information is available.
While the KG-RAG method consistently outperformed the baseline, the experiments also revealed several concrete limitations and failure modes relevant to practical use. The most direct limitation is the dependency on KG coverage. When the KG lacks examples for a specific region of the problem space, such as uncommon facility scales or rare combinations of constraints, the precision of exact-match retrieval decreases, and the system relies more heavily on cluster-based fallbacks. This behaviour was clearly reflected in the sixth test case (125 facilities), where sparse upper-scale coverage caused both methods to depend on coarse cluster-level evidence, resulting in a non-ground-truth recommendation with a lower reasoning score. This reliance ensures that recommendations remain available, but it can reduce specificity at the algorithm–instance level. A second limitation arises from the ingestion process. If user-submitted cases are incorrectly formatted or contain low-quality information, they may introduce noisy or misleading links into the KG, which in turn can affect retrieval accuracy. Third, although grounding substantially reduces the risk of hallucination, it does not fully eliminate overgeneralization. When the assembled dossier for the LLM contains sparse or weak evidence, the model may still generate plausible but insufficiently supported explanations. Fourth, scaling to very large deployments increases computational demands: repeated embedding generation and exhaustive similarity searches can introduce latency and backend load that affect real-time responsiveness. These patterns were evident in the evaluation whenever exact algorithm precedents were missing. In such cases, the system produced cluster-level recommendations that were logically sound but lacked the precision of instance-level matches. Another boundary condition appears in out-of-scope queries, where the problem specifications fall far outside the range of KG data. Rather than failing outright, the system defaults to the most relevant high-level clusters, providing a general but still actionable recommendation. This behaviour illustrates both the resilience and the boundaries of the framework: it can continue to return useful outputs, but high-precision recommendations require representative KG coverage, careful ingestion, and efficient retrieval management.
To address the observed limitations without changing the overall KG-RAG design, several practical measures can be applied. During ingestion, attaching provenance metadata and simple quality scores allows the retrieval engine to weight evidence by reliability, while a lightweight human-in-the-loop review prevents long-term degradation of the KG. Embeddings can be improved by adding structured domain features, such as facility counts or aspect ratios, alongside text descriptions to enhance vector search fidelity. Prompt templates and dossier rules should enforce a minimum evidence threshold before returning strong claims; when this is not met, the system can provide lower-confidence recommendations and suggest data enrichment. Operationally, incremental embedding updates, approximate nearest-neighbour indices, and caching reduce latency as the graph grows. Further analyses would clarify robustness and the contribution of individual components. Ablation studies disabling each retrieval channel (graph-only, vector-only, or cluster-only) would quantify their marginal benefits. Sensitivity analysis masking parts of the KG would reveal how coverage affects performance and guide data acquisition. User studies with domain experts would assess trust and usability, while latency profiling across KG sizes would inform engineering trade-offs. The results also indicate conditions that affect deployment. Performance is strongest when the KG contains diverse, well-labelled precedents covering the target problem range; the behaviour observed in the sixth test case, where sparse upper-scale coverage reduced retrieval precision, highlights the importance of continued data growth. When exact precedents are missing, the fallback to cluster-level evidence produces consistent but less specific recommendations; presenting these with confidence indicators helps practitioners interpret them appropriately. The explicit evidence trail ensures recommendations remain transparent and auditable, supporting the system’s role as a decision-support tool.
@@ 4 Conclusion
Finally, this section concludes by summarizing the primary contributions and key findings of the paper, discussing the current limitations of the method and highlighting promising directions for future research. This paper introduced a new recommendation method for the FLP based on a KG-RAG framework. The method integrates a domain-specific KG with a multi-stage retrieval pipeline and an LLM to deliver explainable, data-driven algorithm recommendations. In evaluation, the proposed method consistently outperformed the baseline Gemini 1.5 Flash LLM provided with the same knowledge base in tabular form, achieving an overall average reasoning score of approximately 4.7 across six test cases, compared to about 3.3 for the baseline. These results highlight the strength of combining structured knowledge with hybrid retrieval strategies, enabling the system to generate context-aware and explainable algorithm recommendations, even in complex scenarios where simpler baseline approaches may struggle. The primary contribution of this work is the design of a multi-stage KG-RAG pipeline that combines graph-based, vector-based, and cluster-based search to assemble context-specific evidence. A key distinction from the baseline is the use of a structured KG rather than a flat CSV table, which enables reasoning over formally encoded relationships instead of relying only on superficial matches. This design leads directly to two major benefits: first, the ability to generate explainable recommendations supported by traceable evidence from problem nodes, method nodes, and cluster trends; and second, a substantial reduction in unsupported or hallucinated outputs, as the LLM is grounded in structured domain knowledge. Together, these contributions demonstrate how the proposed approach advances beyond standard LLM-based recommendations and provides a reliable decision-support framework for FLP.
The study also revealed boundaries that define the current scope of the method. Performance depends on the coverage and quality of the underlying KG, and the sixth test case, which occurs near the upper limit of the current KG illustrated how sparse coverage can reduce retrieval precision and lead to broader, cluster-level recommendations. Recommendations in out-of-scope queries are therefore necessarily based on categorical evidence rather than precise precedents. Ingestion of noisy or low-quality cases may also affect retrieval fidelity, and scaling to very large KG introduces computational demands. These limitations are not fundamental flaws but natural boundaries, and they can be progressively addressed through provenance-aware ingestion, efficient retrieval mechanisms, and continuous graph expansion via the feedback loop. Looking ahead, future work will extend the framework to other combinatorial optimization problems, evaluate scalability across larger and more diverse datasets, and further strengthen continuous update mechanisms. By combining structured knowledge with hybrid retrieval and language model reasoning, the KG-RAG framework establishes a robust foundation for explainable and data-driven algorithm recommendations in complex optimization domains.